OpenStudy (anonymous):

PLEASE HELP! (2x^-3)^3 * (4x^-2)

OpenStudy (anonymous):

first use this formula (x^m)^n = x^(m*n) for (2x^-3)^3

OpenStudy (anonymous):

What do you get ?

OpenStudy (anonymous):

On my first try I got -4x^2 - 32x

OpenStudy (anonymous):

(x^m)^n = x^(m*n) for (2x^-3)^3 So we will get 2x^-9 right ?

OpenStudy (anonymous):

Sorry, I'm a little confused. What does the m and n stand for?

OpenStudy (anonymous):

both are exponents

OpenStudy (anonymous):

what we learned to do in school is first multiply the expiate to the exponent inside the ()

OpenStudy (anonymous):



OpenStudy (anonymous):